OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4)

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Same thing.

OpenStudy (igreen):

\(\sf y - y_1 = m(x - x_1)\) Where \(\sf y_1\) is the y-value of the point, \(\sf x_1\) is the x-value of the point, and \(\sf m\) is the slope.

OpenStudy (anonymous):

I thought the equation for slope form was y=m(x) + b

OpenStudy (anonymous):

so y- -4=-1/3(x-6)

OpenStudy (anonymous):

but then don't the y--4 turn into y+4 @iGreen

OpenStudy (anonymous):

yes it does because two negatives make a positive